Two forces are acting on a body. One acts east, the other at 35° northeast. If the two forces are equal in magnitude of 50 N, find the resultant.
1
Expert's answer
2021-10-28T08:52:16-0400

The resultant is


"R_x=50+50\cos35°=91\text{ N},\\
R_y=50\sin35°=29\text{ N},\\
R=\sqrt{R_x^2+R_y^2} =96\text{ N},\\\space\\
\theta=\arctan\frac{R_y}{R_x}=18°."